OpenStudy (anonymous):

find the slope and the y intercept 3y+2x+9=10+2x

OpenStudy (saifoo.khan):

slope: 0 y intercept = -1/3

OpenStudy (anonymous):

y intercept is not correct

OpenStudy (anonymous):

i got y=+1/3

OpenStudy (saifoo.khan):

my bad, 1/3

myininaya (myininaya):

3y=10+2x-2x-9 3y=1 y=1/3 slope=0 y-interccept=1/3

OpenStudy (anonymous):

3y+2x+9=10+2x 3y=10+2x-2x-9 y = 1/3